# Phase transition like behaviors of Propagation of Passenger Stranding phenomena in Subway Networks

### 12 Jan 2024 v1

@@ Appendix B Supplementary text
B.1 Quantitative definition of cluster
The number of stranded passengers at station $i$ at time $t$ is denoted as $S_{i}(t)$. The relation between two stranded stations, $i$ and $j$, can be classified into two categories:
$e_{ij}=\begin{cases}1&\text{ $S_{i}S_{j}>0$ and $T_{ij}=1$ }\\ 0&\text{ else }\end{cases}$ (B.2)
@@ B.2 System I: Steady-State features
B.2.1 Time Translation Invariance of the Model
When the demand scale is relatively small, such as $P_{u}=20$, the evolution of the three order parameters over time is shown in Fig. B.13a. In the initial stages of the simulation, all vehicles in the system depart from the originating station, while passenger demands are randomly generated throughout the network. Due to insufficient service supply in the early stages, all three order parameters increase rapidly. However, after a relaxation time of $2.0\times 10^{4}$, they reach a steady state. As the demand scale $P_{u}$ increases (as shown in Fig. B.13B for $P_{u}=30$), the system reaches a steady state as long as the relaxation time is sufficiently long. In the steady state, all the order parameters significantly increase with the increasing demand scale $P_{u}$. We use the average values of the 3 order parameters after the system reaches a steady state to characterize the state of the system under a given passenger demand scale $P_{u}$ condition.
Figure B.13: (A) The variation curves of n_congestion, max_cluster, and tot_congestion in the system when $P_{u}=20$. (B) The variation curves of n_congestion, max_cluster, and tot_congestion in the system when $P_{u}=30$. In fact, for all $P_{u}$ values in the system, the four variables eventually reach a steady state. For clarity, only the changes in the number of clusters in the system for two $P_{u}$ values are plotted.
@@ B.2.2 Distribution Evolution of Station with stranded passengers in ideal model
Figure B.14: Evolution of station distribution for the ideal model with $P_{u}=20$.
Figure B.15: Evolution of station distribution for the ideal model with $P_{u}=30$.
@@ B.3.1 Tidal characteristics of demand and supply
Figure B.16: (A) Temporal distribution of demand and supply. Note: $N_{supply}$ represents the total number of trains departing every 10 min in the model shown in Fig. A.11, and $N_{need}$ represents the total number of passengers entering the station every 10 min in the model shown in Fig. A.11. (B) Spatial distribution of passenger demand between 6:00 and 13:00. The diameter of the circles is proportional to the number of passengers entering the station at each location. Larger circles indicate higher passenger volumes, while smaller circles indicate lower volumes.
B.3.2 Analysis of Cluster Evolution
@@ B.3.3 Clusters and Labeling
During the evolution process of the real supply-demand system, if adjacent stations have experienced stranded passengers, they are considered to belong to the same cluster. During the morning peak hours on weekdays, four stranded station clusters have formed in four directions in Beijing (see Fig. B.18). The stations included in each cluster are listed in Table B.6 and Table B.7.
Figure B.17: (A)-(D) Spatial distribution of stranded stations at 7:20, 8:00, 8:20, and 13:00 corresponding to $P_{r}=1.0$. (F)-(I) Spatial distribution of stranded stations at 6:40, 7:30, 8:20, and 13:00 in the network corresponding to $P_{r}=1.6$. (E) and (J) Clusters formed by stations that experienced stranded situations during the morning peak hours (6:30-13:00). These clusters consist of adjacent stations in both spatial and temporal dimensions. The stations included in Cluster I-IV 4 are represented by green, blue, yellow, and purple colors, respectively.
Table B.6: Stations and Station Labels included in the 4 Stranded Station Clusters Formed during Weekday Morning Peak Hours with Passenger Demand of 1.0.
Cluster[c] Station(Label)[c]
@@
Table B.7: Stations and Station labels included in the 4 Clusters formed during weekday Morning rush hour with $P_{r}=1.6$.
Cluster[c] Station(Label)[c]
@@ B.4 Validating the Mechanism of Cluster Formation
B.4.1 Validating the Mechanism of Cluster II-IV
Figure B.18: (A)-(C) Topological Pattern, Nonlocal Pattern, and Local Mode of Cluster II.
Figure B.19: (A)-(C) Topological Pattern, Nonlocal Pattern, and Local Mode of Cluster III.
Figure B.20: (A)-(C) Topological Pattern, Nonlocal Pattern, and Local Mode of Cluster IV.
B.4.2 Train Control Scheme
@@
We maintain the spatial distribution of train departures while adjusting the frequency of train departures on Lines 5 and 13 to modify the supply scale. Initially, we preserve the spatial distribution of train departures according to the original timetable for Lines 5 and 13. We double the train departure density, resulting in a total number of train services denoted as $N_{tot}$. The additional trains in the expanded timetable are scheduled to depart 2 min later than the original timetable (Fig. B.21). Since the time delay is relatively short, the new timetable preserves the spatiotemporal characteristics of the original one. Based on the new train timetable, we randomly select the required number of train services according to passenger demand to form the actual subway timetable. Let $N_{ori}$ denote the total number of train services in the original timetable. We define the adjusted supply, $P_{s}$, as $N_{tot}/N_{ori}$.
@@
Figure B.21: Changes in the Number of Train Services for the Original Timetable ($N_{o}ri$), Additional Timetable ($N_{a}dd$), and New Timetable ($N_{t}ot$) for Lines 5 and 13.
B.4.3 Changes in the Contributions of Different Supply Scale Patterns
Figure B.22: Proportional contributions of Local (green), Nonlocal (red), and Topological (blue) Patterns for different $P_{s}$.

### 28 Feb 2024 v2

@@ Appendix B Supplementary text
B.1 Quantitative definition of the cluster
The number of stranded passengers at station $i$ at time $t$ is denoted as $S_{i}(t)$. The relation between 2 stranded stations, $i$ and $j$, can be classified into 2 categories:
$e_{ij}=\begin{cases}1&\text{ $S_{i}S_{j}>0$ and $T_{ij}=1$ }\\ 0&\text{ else }.\end{cases}$ (12)
@@ B.2 System I: Steady-state features
B.2.1 Temporal variation of the average waiting time and the 3 order parameters
When the demand scale is relatively small, such as $P_{\mathrm{u}}=20$, the evolution of the 3 order parameters over time is shown in Fig. 10a. In the initial stages of the simulation, all vehicles in the system depart from the originating station, while passenger demands are randomly generated throughout the network. Due to insufficient service supply in the early stages, all 3 order parameters increase rapidly. However, after a relaxation time of $2.0\times 10^{4}$, they reach a steady state. As the demand scale $P_{\mathrm{u}}$ increases (as shown in Fig. 10b for $P_{\mathrm{u}}=30$), the system reaches a steady state as long as the relaxation time is sufficiently long. In the steady state, all the order parameters significantly increase with the increasing demand scale $P_{\mathrm{u}}$. We use the average values of the 3 order parameters after the system reaches a steady state to characterize the state of the system under a given passenger demand scale $P_{\mathrm{u}}$ condition.
Figure 10: a The variation curves of $n_{\mathrm{s}}$, $m_{\mathrm{c}}$, and $n_{\mathrm{t}}$ in the system when $P_{\mathrm{u}}=20$. b The variation curves of $n_{\mathrm{s}}$, $m_{\mathrm{c}}$, and $n_{\mathrm{t}}$ in the system when $P_{\mathrm{u}}=30$. In fact, for all $P_{\mathrm{u}}$ values in the system, the 4 variables eventually reach a steady state. For clarity, only the changes in the number of clusters in the system for 2 $P_{\mathrm{u}}$ values are plotted.
@@ B.2.2 Distribution evolution of station with stranded passengers in ideal model
Figure 11: Evolution of station distribution for the ideal model with $P_{\mathrm{u}}=20$.
Figure 12: Evolution of station distribution for the ideal model with $P_{\mathrm{u}}=30$.
@@ B.3.1 Tidal characteristics of demand and supply
Figure 13: a Temporal distribution of demand and supply. Note: $N_{\mathrm{supply}}$ represents the total number of trains departing every 10 min in the model shown in Fig. 8, and $N_{\mathrm{need}}$ represents the total number of passengers entering the station every 10 min in the model shown in Fig. 8. b Spatial distribution of passenger demand between 6:00 and 13:00. The diameter of the loop line is proportional to the number of passengers entering the station at each location. Larger loop lines indicate higher passenger volumes, while smaller loop lines indicate lower volumes.
B.3.2 Dynamics mechanism of cluster growth
@@ B.3.3 Clusters and labeling
During the evolution process of the real supply-demand system, if adjacent stations have experienced stranded passengers, they are considered to belong to the same cluster. During the morning rush hours on weekdays, 4 stranded station clusters have formed in 4 directions in Beijing (Fig. 14 presents critical stations for the 4 clusters for $P_{\mathrm{r}}=1.0$ at $P_{\mathrm{r}}=1.6$). The critical stations included in each cluster are listed in Table 6 and Table 7.
Figure 14: a-d Spatial distribution of critical stations in 4 clusters at 7:20, 8:00, 8:20, and 13:00 corresponding to $P_{\mathrm{r}}=1.0$. f-i Spatial distribution of critical stations in 4 clusters at 6:40, 7:30, 8:20, and 13:00 in the network corresponding to $P_{\mathrm{r}}=1.6$. e and j Clusters formed by critical stations that experienced stranded situations during the morning rush hours (6:30-13:00). These clusters consist of adjacent critical stations in both spatial and temporal dimensions. The stations included in Cluster I-IV 4 are represented by green, blue, yellow, and purple colors, respectively.
Table 6: Critical stations and critical stations’ Label included in the 4 clusters formed during weekday morning rush hours at $P_{\mathrm{r}}=1.0$
Cluster Station(Label)
@@
Table 7: Critical stations and critical stations’ Label included in the 4 clusters formed during weekday morning rush hours at $P_{\mathrm{r}}=1.6$
Cluster Station(Label)
@@ B.4 Validating the mechanism of cluster formation
B.4.1 Dynamics mechanism of cluster I growth
Local pattern: The largest eigenvalue, responsible for 85.9% of the total contribution, exhibits significant spatiotemporal variations at 3 distinct stations: $i=87$, 231, and 232. The temporal evolution of ${{(A}_{1}^{\mathrm{e}})}_{it}$ at these stations is depicted in Fig. 15a, demonstrating a characteristic pattern with a single peak during the early peak period, reaching its maximum around 8:50, and returning to the initial state by approximately 10:40. Station 231 exhibits the most prominent amplitude, while station 87 shows a more gradual variation. These observed patterns coincide with the temporal changes in the number of passengers entering the corresponding stations (i.e., demand scale), suggesting that this stranded pattern arises from an increase in the same passenger demand. Thus, we term this phenomenon the Local pattern. Consequently, the primary driving factor behind the increase in stranded passengers at these stations is the mismatch between supply and demand caused by localized demand growth.
Figure 15: a Temporal evolution of the largest eigenstates in Cluster I for stations 87, 231, and 232 from 6:00 to 13:00. Stations 85, 86, 200, 88, and 89 do not exhibit significant variations in their eigenstates and are therefore not shown. b Variation in the number of entering passengers for stations 87, 231, and 232 from 6:00 to 13:00.
Nonlocal pattern: The second-largest eigenvalue accounts for 13.4% of the total variance. Figs. 16a, c and e depict the temporal evolution of ${{(A}_{2}^{\mathrm{e}})}_{it}$ for stations 87, 231, and 232, respectively. In contrast to the characteristic single peak of Local pattern, this pattern exhibits alternating positive and negative variations, resembling the derivative of a single peak function. By comparing the morning rush hour stranded passenger counts at upstream stations $i=86$, 87, and 231 (Figs. 16b, d, f) with ${{(A}_{2}^{\mathrm{e}})}_{it}$, it can be inferred that the mechanism of stranded conditions involves propagation. This suggests that the variations in this pattern are driven by an increase in boarding passengers at upstream stations during rush hour periods. The stranded conditions at upstream stations result in a reduction of service supply downstream, leading to an increase in stranded passengers at downstream stations. We named this phenomenon the Nonlocal pattern. Specifically, between 7:50 and 9:00, the increase in boarding passengers at station 86 leads to overcrowding, and this overcrowded condition does not improve at the nearest downstream station, 87. This interdependent relationship between adjacent stations facilitates the propagation of stranded conditions along the line direction. This effect is a secondary consequence of the Local pattern, resulting in a lower contribution from the Nonlocal pattern. Similarly, the increase in boarding passengers at station 231 exacerbates the stranded conditions downstream by reducing supply. During this period, both station 87 and station 231 experience stranded conditions, which have a suppressing effect on downstream station 232. Between 9:00 and 10:40, as the number of boarding passengers at station 231 increases, the stranded conditions at downstream station 232 worsen. Simultaneously, the decrease in boarding passengers at station 86 gradually reduces the occupation of supply at downstream stations, thereby alleviating the stranded conditions. Additionally, the impact of station 87 on the stranded conditions at station 231 starts to weaken. This mode reflects the propagation of stranded conditions and explains how the traffic supply at upstream stations influences the stranded conditions at downstream stations.
Figure 16: Variation of the second-largest eigenstates and passenger demand in Cluster I. a, c, and e depict the temporal evolution of the second-largest eigenstates for stations 87, 231, and 232, respectively, from 6:00 to 13:00. b, d, and f represent the changes in the number of boarding passengers at upstream stations 86, 87, and 231, respectively. The eigenstates of other stations within Cluster I exhibit minimal variations.
Topological pattern: The third-largest eigenvalue corresponds to a pattern contribution of 1.7%. Fig. 5 illustrates the temporal evolution of ${{(A}_{3}^{\mathrm{e}})}_{it}$. This pattern unveils the influence of the infrastructure’s topological structure on passenger stranding, which we refer to as the Topological pattern. The distinguishing characteristic of this pattern is the significant variation observed at a specific station, designated as Local pattern, as depicted in Fig. 17 for station 87. Unlike the other 2 stations, station 87 functions as a transfer station, with multiple upstream stations contributing to the propagation of stranding and multiple downstream stations contributing to stranding alleviation, while stations 231 and 232 serve as non-transfer stations. Consistent with the findings from steady-state condition calculations under spatially and temporally homogeneous demand conditions discussed in the previous section, stations with high network betweenness centrality at transfer points play a pivotal role in PPS occurring. Under steady-state conditions, only the topological structure of the network influences PPS. In this system characterized by spatially and temporally heterogeneous supply and demand, the contribution of the topological factor ranks third.
B.4.2 Dynamics mechanism of II-IV growth
Figure 18: a-c Topological Pattern, Nonlocal Pattern, and Local Mode of Cluster II.
Figure 19: a-c Topological Pattern, Nonlocal Pattern, and Local Mode of Cluster III.
Figure 20: a-c Topological pattern, Nonlocal pattern, and Local pattern of Cluster IV.
B.4.3 Train control scheme
@@
We maintain the spatial distribution of train departures while adjusting the frequency of train departures on Lines 5 and 13 to modify the supply scale. Initially, we preserve the spatial distribution of train departures according to the original timetable for Lines 5 and 13. We double the train departure density, resulting in a total number of train services denoted as $N_{\mathrm{tot}}$. The additional trains in the expanded timetable are scheduled to depart 2 min later than the original timetable (Fig. 21). Since the time delay is relatively short, the new timetable preserves the spatiotemporal characteristics of the original one. Based on the new train timetable, we randomly select the required number of train services according to passenger demand to form the actual subway timetable. Let $N_{\mathrm{ori}}$ denote the total number of train services in the original timetable. We define the adjusted supply, $P_{\mathrm{s}}$, as $N_{\mathrm{tot}}/N_{\mathrm{ori}}$.
@@
Figure 21: Changes in the Number of Train Services for the Original Timetable ($N_{\mathrm{ori}}$), Additional Timetable ($N_{\mathrm{add}}$), and New Timetable ($N_{\mathrm{tot}}$) for Lines 5 and 13.
B.4.4 Changes in the contributions of different supply scale patterns
[…]